• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name hoyoung

    Log entry time 00:01:00 on October13,2009

    Entry number 294637

    keyword=Swing shift summary

    Shift workers: Hoyoung Kang(SL), Ge Jin(TO), Chiranjib Dutta(3rd)

    16:00 During 4-5pm, beam current reached at 80uA for 10 minutes and somtimes 95uA. But they were not acceptable. We took pedestal runs.
    17:40 Beam is stable. Harp scan was requested with raster off.
    18:20 Beam position correction. Target moved to carbon and We did spot++ at 10uA.
    18:30 LH2 target. spot++ at 10uA.
    18:37 Started production run 13209 at 100uA.
    20:53 Routine spot++ during production
    22:36 CODA reboot
    23:00 Routine spot++ during production

    Until run# 13214, we have archived 314.2k OK pairs with IHWP-IN.

    13207 pedestal
    13208 pedestal
    13209 production 100uA(occasional beam trip during first 10 min) 5.8x5
    13210 production 100uA 5.8x5
    13211 production 100uA 5.8x5
    13212 production 100uA 5.8x5
    13213 short run(coda crash)
    13214 production 100uA 5.8x5
    13215 production 100uA 5.8x5